# MLaKE: Multilingual Knowledge Editing Benchmark for Large Language Models

### v1

@@ 2.2 Description of the MLaKE
Dataset statistics
Table 1 summarizes the statistics of the MLaKE dataset. The MLaKE dataset consists of more than 13K samples, each of which has both free-question (Free-form QA) and multiple-choice (MCQA) formats. We align all single-hop problems and employ them as a means to investigate the generalizability of existing knowledge editing methods across different languages following editing in a single language. Multi-hop problems are not aligned across different languages, and we employ them to assess the portability of existing editing methods. Each multi-hop problem (original) is linked to a real-world fact. Our dataset encompasses a broad spectrum of relationships and entities, showcasing a rich diversity within its contents.
Dataset analysis
Figure 2 briefly analyzes the characteristics of the MLaKE dataset. Figure 2(a) depicts all first relations and their corresponding top 3 second relations, demonstrating the diversity of relations in MLaKE. The majority of questions pertain to nationality, names of individuals and locations, and typically adhere to the following structure: "From which nation does Gwendoline Christie hold citizenship?" (single-hop question) or "In which country is the institution where Carl Sagan was employed located?" (multi-hop question). In Figure 2(b), we further examine the relation PIDs that account for more than 1%, mainly including "country of citizenship" (p27), "country" (P17), "continent" (P30), etc. For the corresponding table of relationship PID and relationship label, please refer to the Appendix A. Figure 2(c) depicts the distribution of entities that have an occurrence rate exceeding 0.5%. The prominent entities in this distribution include ’United Kingdom’ (Q145), ’Canada’ (Q36), and ’United States of America’ (Q30). Figure 2(d) illustrates the distribution of question lengths. The majority of questions fall within the 10-20 word range, which allows for precise expression of the subject and relationship without the inclusion of extraneous information. To accommodate various answer preferences of Large Language Models (LLMs), we strive to generate multiple aliases for each answer in MLaKE. Figure 2(e) demonstrates that the majority of answers possess 2-13 aliases, while there are even several answers with more than 20 aliases. The analysis data presented in Figure 2 only includes English samples, both single-hop and multi-hop.
@@
Refer to caption (b) Frequency distribution of relations (only occurrences $>$1% are shown). Refer to caption (c) Frequency distribution of entities (only occurrences $>$0.5% are shown).
Refer to caption (d) Refer to caption (e)
@@
2.3 Evaluation Metrics
Diverging from other benchmarks, MLaKE primarily focuses on assessing the transferability of knowledge editing methods across multilingual scenarios. For different models, we use corresponding prompt templates to guide them to generate answers. The evaluation criteria for different tasks are as follows: 1) QA Accuracy: We detect the presence of edited answers or their aliases in the sentences generated by the models. 2) MCQA Accuracy: We calculate the probabilities of the tokens "A", "B", "C" and "D" generated by the model and take the option with the highest probability as the prediction.
3 Experiments
In this section, we first introduce the experimental setups in our study, including foundation models, knowledge editing baselines and implementation details, and then carefully analyze the experimental results.
3.1 Experimental Setup
Language Models
Considering the necessity for multilingual capabilities, we incorporate the following five language models in our experiments: 1) Llama-2-7B-Chat is optimized based on the Llama-2 model (Touvron et al., 2023) specifically for dialogue-oriented applications through supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF) to align with human preferences regarding helpfulness and safety. 2) Vicuna-7B-v1.5 is fine-tuned from Llama-2 using supervised instruction fine-tuning with training data sourced from ShareGPT.com. 3) Qwen1.5-7B-Chat is a transformer-based aligned chat model pre-trained on extensive data, developed by Alibaba Cloud. 4) Gemma-7B-IT is an instruction-tuned model employing similar architectures, datasets, and training methodologies as the Gemini model (Team et al., 2024). 5) Mistral-7B-Instruct-v0.2 is an enhanced instruct fine-tuned iteration of the Mistral-7B model (Jiang et al., 2023).
Knowledge Editing Methods
Implementation Details
All experiments are conducted on a single NVIDIA A800 GPU (80GB). We re-implement MEND, ROME and MEMIT using EasyEdit [^3] https://github.com/zjunlp/EasyEdit (Wang et al., 2023c) with default settings. Additionally, we reproduce the results of StableKE by utilizing their official repository [^4] https://github.com/Hi-archers/StableKE. In the case of MEND, MEMIT, and StableKE, a batch size of 100 was employed for actual editing. As for ROME, which does not support batch editing, a batch size of 1 was used, and 100 samples were iteratively edited before testing. See the code for detailed settings.
Models Single-hop Free-form QA (%) Single-hop Multi-Choice QA (%)
EN ZH JA FR DE EN ZH JA FR DE
LLaMa-2-7B-Chat 78.17 0.0 0.0 63.62 59.05 85.26 51.40 60.54 79.01 78.64
Vicuna-7B-v1.5 76.21 36.94 46.83 61.29 58.40 85.54 44.59 54.38 82.28 81.16
Qwen1.5-7B-Chat 78.92 57.00 37.87 58.21 54.85 89.83 76.59 64.55 88.53 87.78
Gemma-7B-IT 50.37 34.33 43.00 42.35 42.07 80.97 48.97 55.60 70.99 72.29
Mistral-7B-Ins.-v0.2 83.21 28.17 26.77 65.95 60.73 88.99 44.22 45.43 87.69 85.82
Table 2: Single-hop question answering performance of five LLMs on MLaKE. Based on our experiments, it was observed that LLaMa-2-7B-Chat lacks the capability to generate responses in Chinese and Japanese, resulting in a QA result of 0.
Models Multi-hop Free-form QA (%) Multi-hop Multi-Choice QA (%)
EN ZH JA FR DE EN ZH JA FR DE
LLaMa-2-7B-Chat 49.94 0.0 0.0 29.92 25.99 60.81 44.74 47.23 63.81 55.16
Vicuna-7B-v1.5 47.60 21.32 26.50 26.98 28.50 62.77 34.74 43.46 65.09 55.29
Qwen1.5-7B-Chat 60.81 53.55 30.51 30.56 30.46 58.62 58.29 43.35 68.41 57.91
Gemma-7B-IT 34.83 22.11 24.26 22.51 12.94 56.33 44.61 45.35 53.45 48.76
Mistral-7B-Ins.-v0.2 62.12 29.21 18.85 37.21 40.39 64.85 43.03 40.52 69.95 63.01
Table 3: Multi-hop question answering performance of five LLMs on MLaKE.
3.2 Experimental Results
3.2.1 Evaluating QA Performance of Five Widely Used LLMs on MLaKE
To explore the extent of knowledge acquisition in each language by LLMs, we first evaluate the question answering performance of five widely used LLMs on MLaKE, and the experimental results of single-hop and multi-hop QA are shown in Table 2 and 3, respectively. Notably, LLaMA-7B-Chat lacks the capability to generate responses in Chinese and Japanese, hence its performance on free-form QA in these languages is unavailable. Our key observations include: (1) there are significant discrepancies in the QA performance across different languages, even for the same model. Specifically, English consistently demonstrates higher accuracy across both QA settings, suggesting that LMs tend to exhibit a more comprehensive grasp of knowledge in English. (2) Qwen1.5-7B-Chat displays a more balanced QA performance across various languages compared to other LLMs, particularly excelling in Chinese and Japanese, as its pre-training corpus covers a wide range of languages Bai et al. (2023). (3) A comparison between the performances of LLaMa-2-7B-Chat and Vicuna-7B-v1.5 reveals a remarkable similarity. This similarity can be attributed to the fact that both Vicuna-7B-v1.5 and LLaMa-2-7B-Chat are fine-tuned versions of LLaMa-2-7B.
We also note a significant discrepancy in the free-form QA accuracy of Gemma-7B-IT compared to other models. Specifically, it falls behind LLaMa-2-7B-Chat by 27.8% and 15.11% on single-hop and multi-hop QA, respectively. This disparity can be primarily attributed to the frequent occurrence of rejection responses acquired through Gemma’s RLHF process.
Models Single-hop Free-form QA (%) Single-hop Multi-Choice QA (%)
EN ZH JA FR DE EN ZH JA FR DE
MEND 0.09 0.00 0.00 0.00 0.19 19.12 24.81 24.53 11.75 14.46
ROME 14.93 0.00 0.09 3.17 5.50 25.37 24.72 23.32 24.53 24.81
MEMIT 60.73 7.18 9.33 22.85 24.53 35.07 23.04 22.20 32.18 34.79
StableKE 88.43 5.50 8.12 34.14 38.15 23.51 21.64 17.16 22.85 22.67
Table 4: Single-hop knowledge editing performance of existing methods (backbone model Vicuna-7B-v1.5) on MLaKE. For all methods, we edit the model in English but ask it to produce accurate answer with other languages. The language identifier in this table indicates the target language.
Models Multi-hop Free-form QA (%) Multi-hop Multi-Choice QA (%)
EN ZH JA FR DE EN ZH JA FR DE
MEND 0.20 0.25 0.22 0.43 0.55 25.78 26.62 26.51 28.25 28.45
ROME 2.44 0.00 0.11 1.43 2.33 23.44 26.74 25.96 27.25 26.95
MEMIT 20.70 9.13 10.01 14.12 15.18 34.47 29.02 29.95 39.51 42.27
StableKE 26.67 6.21 8.69 15.41 17.10 34.67 33.33 31.02 39.66 44.05
Table 5: Multi-hop knowledge editing performance of existing methods (backbone model Vicuna-7B-v1.5) on MLaKE. The same as the single-hop setting, we also edit in English but test the QA accuracy in other languages.
3.2.2 Knowledge Editing Performance of Existing KE Methods on MLaKE
To evaluate the effectiveness of current knowledge editing methods in multilingual scenarios, we conducted experiments to assess their ability to edit knowledge in the source language (e.g., English) while producing accurate responses in a different target language. Analysis of the results presented in Table 4, Table 5 and Figure 3 reveals the following observations:
  1. (1)
    Existing knowledge editing methods encounter challenges in maintaining accuracy during the transition from the source language to a different target language. Table 4 demonstrates that in single-hop QA, existing methods experience a significant decline in their free-form QA performance after editing in English. This decline results in a near-complete loss of their ability to generate accurate answers in other languages. Similar trends are observed in the results of MCQA. Under our settings, certain knowledge editing methods, such as MEND and ROME, exhibit parameter values that are on the verge of collapse. Consequently, we consider their results as non-informative, as a performance of 25 indicates random guessing. In addition, multi-language and multi-hop knowledge editing poses greater challenges. Table 5 demonstrates that even the most powerful performer in cross-language knowledge editing, StableKE, experiences a significant decrease in editing accuracy compared to its single-hop performance.
  2. (2)
    Existing knowledge editing methods exhibit significantly higher editing accuracy on English samples compared to other languages. Referring to Table 4 and Table 5, it is evident that, except for the knowledge editing method with parameter collapse, the EN metric of the remaining methods outperforms other indicators in the QA task. Moreover, in Figure 3, both MEMIT and StableKE exhibit the highest editing success rate on EN-EN along the diagonal line.
  3. (3)
    In multilingual knowledge editing, existing knowledge editing methods have higher generalization in languages similar to the editing language. A distinct contrast in response accuracy is observed between Chinese and Japanese compared to French and German when editing is conducted in English. This difference is attributed to the greater linguistic divergence between English and Chinese/Japanese as opposed to French/German. To further validate this hypothesis, we expand our experiments to include editing in other languages and subsequently evaluate the accuracy of responses in English. The detailed experimental results in Figure 3 consistently reaffirm the initial observation: the accuracy of responding in English after editing in Chinese and Japanese is considerably lower compared to editing in French and German. This underscores the significant impact of language differences on cross-language knowledge editing performance.
Our findings underscore the complexities inherent in cross-language knowledge editing and emphasize the need for further advancements to achieve more reliable and accurate results.
Figure 3: The capability of StableKE and MEMIT to edit knowledge in the source language and generate accurate responses in a different target language.
4 Related Work
4.1 Knowledge Editing Datasets
4.2 Knowledge Editing Methods
5 Conclusion and Future Work
In this paper, we explore the effectiveness of current knowledge editing methods in multilingual settings. To this end, we created the MLaKE dataset by extracting knowledge tuples from Wikipedia and generating single-hop and multi-hop questions using ChatGPT. Leveraging MLaKE, we conducted experiments employing various methods and multilingual LLMs to investigate the transferability of knowledge editing from English to other languages and vice versa. Our findings suggest that: (1) Language discrepancies among different languages significantly influence the efficacy of knowledge editing. (2) Existing methods still face challenges in editing knowledge in the source language and generating accurate responses in a different language. (3) While LLMs demonstrate the capability to align cross-language knowledge to some extent, conveying this alignment in natural language remains problematic. In the future, we aim to develop reliable multilingual knowledge editing techniques to enhance the cross-language knowledge alignment and expression capabilities of LLMs.
Appendix A Data Details
A.1 Filter rules
Four professionals were employed to meticulously perform data cleaning, systematically addressing any flaws that may have been overlooked. In addition, the filtering data rules we set are as follows:
  1. (1)
@@ A.1 Filter rules
  2. (2)
    Among all data, the relationship corresponding to an entity cannot be modified repeatedly. To facilitate evaluation and avoid knowledge conflicts, we refrain from repeatedly modifying the fact chain of an entity among all the data. This is particularly relevant considering that several knowledge editing methods support batch editing capabilities.
  3. (3)
    The data is devoid of any toxic information pertaining to politics, violence, or pornography.
  4. (4)
    Fact chains, especially multi-hop ones, do not contain any circular dependencies.
  5. (5)
    All chains of facts naturally give rise to corresponding questions.
A.2 Instructions
@@
Question Generation Instruction
Given the Wikidata knowledge triplet structure [subject, relation, object] where subject is $s_{1}$, relation is $r_{1}$, use this information to guide ChatGPT in creating a question that aims to identify the $o_{1}$ of the triplet. Your challenge is to create a detailed question that prompts LLM to identify the $r_{1}$ of $s_{1}$, without giving away the $o_{1}$ or making any reference to it.
Answer Generation Instruction
Using the provided Wikidata knowledge triple [$s_{1}$, $r_{1}$, $o_{1}$], craft a concise answer to the question (question). Your response should clearly link the $s_{1}$ with the $o_{1}$ as the answer, without delving into additional details or context. The aim is to directly address the query with the information given in the triple.
@@
Question Generation Instruction
给定Wikidata知识三元组[subject, relation, object],subject是$s_{1}$,relation是$r_{1}$,。根据这些信息,设计一个中文问题,旨在根据$s_{1}$$r_{1}$询问三元组中的object,即$o_{1}$。问题应详细且具体,让便让LLM准确的回答出$s_{1}$$r_{1}$,同时避免直接提到或暗示$o_{1}$
Answer Generation Instruction
根据给定的Wikidata知识三元组[$s_{1}$, $r_{1}$, $o_{1}$], 就问题(question)形成一个简洁的回答。您的回答应明确地将$s_{1}$$o_{1}$联系起来作为答案,避免深入其他细节或背景。请力求简明扼要,利用特定的三元组关系,确保回答简洁且直接相关。请直接生成结果,不要说无关的内容。
@@
Question Generation Instruction
Wikidataの知識トリプレット構造[subject, relation, object]が与えられた場合、ここでsubjectは$s_{1}$、relationは$r_{1}$です。この情報を使用して、トリプレットの$o_{1}$を特定することを目的とした質問をChatGPTが作成するよう導いてください。あなたの挑戦は、$o_{1}$明かしたり、それに言及したりすることなく、$s_{1}$$r_{1}$を特定するようLLMに促す詳細な質問を作成することです。
Answer Generation Instruction
ウィキデータの知識トリプル [$s_{1}$, $r_{1}$, $o_{1}$] を考慮すると、 、質問「(question)」に対する正確な応答を作成します。あなたの答えは、「$r_{1}$」によって提供されるコンテキストを通じて、「$s_{1}$」を「$r_{1}$」に直接関連付ける必要があります。トリプルで示される特定の関係を活用した簡単な説明を目指し、応答が簡潔で質問に直接関連していることを確認します。
@@
Question Generation Instruction
Étant donné la structure de triplet de connaissances Wikidata [subject, relation, object] où subject est $s_{1}$, relation est $r_{1}$, utilisez cette information pour créer une question avec ChatGPT pour identifier l’élément $o_{1}$ du triplet. Votre défi est de créer une question détaillée qui incite LLM à identifier le $r_{1}$ de $o_{1}$, sans révéler le $o_{1}$ ou faire référence à celui-ci.
Answer Generation Instruction
Étant donné le triplet de connaissances Wikidata [$s_{1}$, $r_{1}$, $o_{1}$], formulez une réponse concise à la question (question). Votre réponse devrait clairement lier le $s_{1}$ avec le $o_{1}$ comme réponse, sans entrer dans des détails ou contextes supplémentaires. Visez une explication simple qui tire parti de la relation spécifique dénotée par le triplet, en vous assurant que la réponse est succincte et directement pertinente à la question. Veuillez générer les résultats directement et ne dites pas de contenu hors sujet.
@@
Question Generation Instruction
Angesichts der Struktur eines Wissens-Tripels in Wikidata (subject, relation, object), wobei ’subject’ $s_{1}$ ist, ’relation’ $r_{1}$ ist, nutzen Sie diese Informationen, um ChatGPT bei der Erstellung einer Frage zu leiten, die darauf abzielt, das $o_{1}$ des Tripels zu identifizieren. Ihre Herausforderung besteht darin, eine detaillierte Frage zu formulieren, die LLM dazu anregt, das $r_{1}$ von $s_{1}$ zu identifizieren, ohne das $o_{1}$ preiszugeben oder darauf Bezug zu nehmen.
Answer Generation Instruction
Verwendung des bereitgestellten Wikidata-Wissenstripels ($s_{1}$, $r_{1}$, $o_{1}$), verfassen Sie eine prägnante Antwort auf die Frage (question). Ihre Antwort sollte $s_{1}$ eindeutig mit $o_{1}$ als Antwort verknüpfen, ohne auf zusätzliche Details oder den Kontext einzugehen. Ziel ist es, mit den im Tripel enthaltenen Informationen direkt auf die Anfrage einzugehen.
@@
Question Generation Instruction
Given the Wikidata triples: ($s_{1}$, $r_{1}$, $o_{1}$) and ($o_{1}$, $r_{2}$, x2), craft a multi-hop question in natural English about $s_{1}$ that explicitly involves the relationships $r_{1}$ and $r_{2}$. The question must ensure there is no implicit or explicit reference to or information leakage about $s_{1}$, leading to an inquiry about x2 without revealing $s_{1}$.
Option Generation Instruction
Given the information: ’question’ and the known facts: ($s_{1}$, $r_{1}$, $o_{1}$) and ($o_{1}$, $r_{2}$, x2), please generate a correct option A and provide three incorrect but plausible options B, C, and D. Ensure that all options are presented in a sentence, not just single words or phrases. The incorrect options should be related enough to the correct answer to pose a challenge, but there’s no need to mention the intermediary connecting entity ($s_{1}$) or any other detailed information.
@@
Question Generation Instruction
请根据以下Wikidata知识三元组:($s_{1}$, $r_{1}$, $o_{1}$)和($o_{1}$, $r_{2}$, x2),用流畅的中文提出一个关于$s_{1}$的多跳问题,用于通过$s_{1}$查询得到x2,该问题必须明确涉及关系$r_{1}$$r_{2}$,从而引出关于x2的询问。(问题要确保没有对$s_{1}$的直接或间接引用或信息泄露)。
Option Generation Instruction
给定信息:’question’,已知事实包括:($s_{1}$, $r_{1}$, $o_{1}$)和($o_{1}$, $r_{2}$, x2)。请根据这些信息,生成一个正确的选项A,并提供三个错误但听起来合理的选项B、C和D。请确保这些选项是以完整的句子形式提出的,而不仅仅是单个词或短语,错误选项应该与正确答案有一定的关联度,使问题具有一定的挑战性,但无需提及中间的连接实体($s_{1}$)或其他详细信息。
@@
Question Generation Instruction
次のWikidataの知識トリプルに基づいて、$s_{1}$についての1つの質問を作成してください:($s_{1}$, $r_{1}$, $o_{1}$)および($o_{1}$, $r_{2}$, x2)。この質問は、$r_{1}$$r_{2}$の関係を通じて$s_{1}$からx2への論理的な連鎖を明確にし、$s_{1}$に関する直接または間接的な参照を避けなければなりません。
Option Generation Instruction
情報「question」と既知の事実「$s_{1}$, $r_{1}$, $o_{1}$」および「$o_{1}$, $r_{2}$, x2」に基づき、正しい選択肢Aを生成し、不正解ながらも妥当な選択肢B、C、Dを3つ提供してください。すべての選択肢は、単語やフレーズだけでなく、文章で提示する必要があります。不正解の選択肢は、正解と十分関連していて挑戦となる必要がありますが、中間の接続エンティティ「$s_{1}$」や他の詳細な情報に言及する必要はありません。
@@
Question Generation Instruction
Veuillez poser une question multi-sauts en français fluide basée sur les triplets de connaissances Wikidata suivants : ($s_{1}$, $r_{1}$, $o_{1}$) et ($o_{1}$, $r_{2}$, x2). La question doit concerner $s_{1}$ et servir à interroger pour obtenir x2, en mentionnant explicitement les relations $r_{1}$ et $r_{2}$ pour introduire une question sur x2. (La question doit s’assurer qu’il n’y ait aucune référence directe ou indirecte à $s_{1}$ ou de fuite d’informations à son sujet).
Option Generation Instruction
Informations fournies: ’question’ Les faits connus incluent: ($s_{1}$, $r_{1}$, $o_{1}$) et ($o_{1}$, $r_{2}$, x2). Sur la base de ces informations, veuillez générer une option correcte A et trois options incorrectes mais raisonnables B, C et D. Assurez-vous que les options sont présentées sous forme de phrases complètes et pas seulement de mots ou d’expressions simples. Les mauvaises options doivent avoir un certain degré de pertinence par rapport à la bonne réponse pour rendre la question difficile, mais sans mentionner l’entité de connexion entre les deux ($s_{1}$) ou d’autres détails.
@@
Question Generation Instruction
Ich möchte eine deutsche Multi-Hop-Frage formulieren, basierend auf den gegebenen Wikidata-Wissens-Tripeln: ($s_{1}$, $r_{1}$, $o_{1}$) und ($o_{1}$, $r_{2}$, x2). Die Frage soll flüssig in deutscher Sprache gestellt werden und es ermöglichen, durch Abfrage von x0 die Information über x2 zu erhalten. Dabei muss die Frage die Beziehungen r1 und r2 explizit einbeziehen, ohne jedoch direkte oder indirekte Hinweise auf die Brückenentität x1 zu geben.
Option Generation Instruction
Gegebene Informationen: ’question’ Zu den bekannten Fakten gehören: ($s_{1}$, $r_{1}$, $o_{1}$) und ($o_{1}$, $r_{2}$, x2). Generieren Sie auf der Grundlage dieser Informationen bitte eine korrekte Option A und drei falsche, aber vernünftig klingende Optionen B, C und D. Stellen Sie sicher, dass die Optionen als vollständige Sätze und nicht nur als einzelne Wörter oder Phrasen dargestellt werden. Die falschen Optionen sollten einen gewissen Grad an Relevanz für die richtige Antwort haben, um die Frage herausfordernd zu gestalten, ohne jedoch die verbindende Entität dazwischen zu erwähnen ($s_{1}$) oder andere Details.
@@
Appendix B Multilingual Knowledge Editing Experiment

### v2

@@ Appendix C Data Details
(b) Frequency distribution of relations (only occurrences $>$1% are shown).
(c) Frequency distribution of entities (only occurrences $>$0.5% are shown).
@@
C.1 Filter rules
We hired five professionals, each specializing in one of the five languages used in this paper, to conduct a comprehensive data cleaning exercise on the dataset according to the following guidelines:
  1. 1.
@@ C.1 Filter rules
  2. 2.
    Among all data, the relationship corresponding to an entity cannot be modified repeatedly. To facilitate evaluation and avoid knowledge conflicts, we ensure that the fact chain of an entity is not modified multiple times. This is particularly relevant considering that several knowledge editing methods support batch editing capabilities.
  3. 3.
    The dataset is free of toxic information, including content related to politics, violence, or pornography.
  4. 4.
    Fact chains, particularly multi-hop ones, are free from circular dependencies.
  5. 5.
    Identify and eliminate any semantic or syntactic errors in the questions and answers generated by the model.
C.2 Instructions
@@
Question Generation Instruction Given the Wikidata knowledge triplet structure [subject, relation, object] where subject is $s_{1}$, relation is $r_{1}$, use this information to guide ChatGPT in creating a question that aims to identify the $o_{1}$ of the triplet. Your challenge is to create a detailed question that prompts LLM to identify the $r_{1}$ of $s_{1}$, without giving away the $o_{1}$ or making any reference to it.
Answer Generation Instruction Using the provided Wikidata knowledge triple [$s_{1}$, $r_{1}$, $o_{1}$], craft a concise answer to the question (question). Your response should clearly link the $s_{1}$ with the $o_{1}$ as the answer, without delving into additional details or context. The aim is to directly address the query with the information given in the triple.
@@
Question Generation Instruction 给定Wikidata知识三元组[subject, relation, object],subject是$s_{1}$,relation是$r_{1}$,。根据这些信息,设计一个中文问题,旨在根据$s_{1}$$r_{1}$询问三元组中的object,即$o_{1}$。问题应详细且具体,让便让LLM准确的回答出$s_{1}$$r_{1}$,同时避免直接提到或暗示$o_{1}$
Answer Generation Instruction 根据给定的Wikidata知识三元组[$s_{1}$, $r_{1}$, $o_{1}$], 就问题(question)形成一个简洁的回答。您的回答应明确地将$s_{1}$$o_{1}$联系起来作为答案,避免深入其他细节或背景。请力求简明扼要,利用特定的三元组关系,确保回答简洁且直接相关。请直接生成结果,不要说无关的内容。
@@
Question Generation Instruction Wikidataの知識トリプレット構造[subject, relation, object]が与えられた場合、ここでsubjectは$s_{1}$、relationは$r_{1}$です。この情報を使用して、トリプレットの$o_{1}$を特定することを目的とした質問をChatGPTが作成するよう導いてください。あなたの挑戦は、$o_{1}$明かしたり、それに言及したりすることなく、$s_{1}$$r_{1}$を特定するようLLMに促す詳細な質問を作成することです。
Answer Generation Instruction ウィキデータの知識トリプル [$s_{1}$, $r_{1}$, $o_{1}$] を考慮すると、 、質問「(question)」に対する正確な応答を作成します。あなたの答えは、「$r_{1}$」によって提供されるコンテキストを通じて、「$s_{1}$」を「$r_{1}$」に直接関連付ける必要があります。トリプルで示される特定の関係を活用した簡単な説明を目指し、応答が簡潔で質問に直接関連していることを確認します。
@@
Question Generation Instruction Étant donné la structure de triplet de connaissances Wikidata [subject, relation, object] où subject est $s_{1}$, relation est $r_{1}$, utilisez cette information pour créer une question avec ChatGPT pour identifier l’élément $o_{1}$ du triplet. Votre défi est de créer une question détaillée qui incite LLM à identifier le $r_{1}$ de $o_{1}$, sans révéler le $o_{1}$ ou faire référence à celui-ci.
Answer Generation Instruction Étant donné le triplet de connaissances Wikidata [$s_{1}$, $r_{1}$, $o_{1}$], formulez une réponse concise à la question (question). Votre réponse devrait clairement lier le $s_{1}$ avec le $o_{1}$ comme réponse, sans entrer dans des détails ou contextes supplémentaires. Visez une explication simple qui tire parti de la relation spécifique dénotée par le triplet, en vous assurant que la réponse est succincte et directement pertinente à la question. Veuillez générer les résultats directement et ne dites pas de contenu hors sujet.
@@
Question Generation Instruction Angesichts der Struktur eines Wissens-Tripels in Wikidata (subject, relation, object), wobei ’subject’ $s_{1}$ ist, ’relation’ $r_{1}$ ist, nutzen Sie diese Informationen, um ChatGPT bei der Erstellung einer Frage zu leiten, die darauf abzielt, das $o_{1}$ des Tripels zu identifizieren. Ihre Herausforderung besteht darin, eine detaillierte Frage zu formulieren, die LLM dazu anregt, das $r_{1}$ von $s_{1}$ zu identifizieren, ohne das $o_{1}$ preiszugeben oder darauf Bezug zu nehmen.
Answer Generation Instruction Verwendung des bereitgestellten Wikidata-Wissenstripels ($s_{1}$, $r_{1}$, $o_{1}$), verfassen Sie eine prägnante Antwort auf die Frage (question). Ihre Antwort sollte $s_{1}$ eindeutig mit $o_{1}$ als Antwort verknüpfen, ohne auf zusätzliche Details oder den Kontext einzugehen. Ziel ist es, mit den im Tripel enthaltenen Informationen direkt auf die Anfrage einzugehen.
@@
Question Generation Instruction Given the Wikidata triples: ($s_{1}$, $r_{1}$, $o_{1}$) and ($o_{1}$, $r_{2}$, x2), craft a multi-hop question in natural English about $s_{1}$ that explicitly involves the relationships $r_{1}$ and $r_{2}$. The question must ensure there is no implicit or explicit reference to or information leakage about $s_{1}$, leading to an inquiry about x2 without revealing $s_{1}$.
Option Generation Instruction Given the information: ’question’ and the known facts: ($s_{1}$, $r_{1}$, $o_{1}$) and ($o_{1}$, $r_{2}$, x2), please generate a correct option A and provide three incorrect but plausible options B, C, and D. Ensure that all options are presented in a sentence, not just single words or phrases. The incorrect options should be related enough to the correct answer to pose a challenge, but there’s no need to mention the intermediary connecting entity ($s_{1}$) or any other detailed information.
@@
Question Generation Instruction 请根据以下Wikidata知识三元组:($s_{1}$, $r_{1}$, $o_{1}$)和($o_{1}$, $r_{2}$, x2),用流畅的中文提出一个关于$s_{1}$的多跳问题,用于通过$s_{1}$查询得到x2,该问题必须明确涉及关系$r_{1}$$r_{2}$,从而引出关于x2的询问。(问题要确保没有对$s_{1}$的直接或间接引用或信息泄露)。
Option Generation Instruction 给定信息:’question’,已知事实包括:($s_{1}$, $r_{1}$, $o_{1}$)和($o_{1}$, $r_{2}$, x2)。请根据这些信息,生成一个正确的选项A,并提供三个错误但听起来合理的选项B、C和D。请确保这些选项是以完整的句子形式提出的,而不仅仅是单个词或短语,错误选项应该与正确答案有一定的关联度,使问题具有一定的挑战性,但无需提及中间的连接实体($s_{1}$)或其他详细信息。
@@
Question Generation Instruction 次のWikidataの知識トリプルに基づいて、$s_{1}$についての1つの質問を作成してください:($s_{1}$, $r_{1}$, $o_{1}$)および($o_{1}$, $r_{2}$, x2)。この質問は、$r_{1}$$r_{2}$の関係を通じて$s_{1}$からx2への論理的な連鎖を明確にし、$s_{1}$に関する直接または間接的な参照を避けなければなりません。
Option Generation Instruction 情報「question」と既知の事実「$s_{1}$, $r_{1}$, $o_{1}$」および「$o_{1}$, $r_{2}$, x2」に基づき、正しい選択肢Aを生成し、不正解ながらも妥当な選択肢B、C、Dを3つ提供してください。すべての選択肢は、単語やフレーズだけでなく、文章で提示する必要があります。不正解の選択肢は、正解と十分関連していて挑戦となる必要がありますが、中間の接続エンティティ「$s_{1}$」や他の詳細な情報に言及する必要はありません。
@@
Question Generation Instruction Veuillez poser une question multi-sauts en français fluide basée sur les triplets de connaissances Wikidata suivants : ($s_{1}$, $r_{1}$, $o_{1}$) et ($o_{1}$, $r_{2}$, x2). La question doit concerner $s_{1}$ et servir à interroger pour obtenir x2, en mentionnant explicitement les relations $r_{1}$ et $r_{2}$ pour introduire une question sur x2. (La question doit s’assurer qu’il n’y ait aucune référence directe ou indirecte à $s_{1}$ ou de fuite d’informations à son sujet).
Option Generation Instruction Informations fournies: ’question’ Les faits connus incluent: ($s_{1}$, $r_{1}$, $o_{1}$) et ($o_{1}$, $r_{2}$, x2). Sur la base de ces informations, veuillez générer une option correcte A et trois options incorrectes mais raisonnables B, C et D. Assurez-vous que les options sont présentées sous forme de phrases complètes et pas seulement de mots ou d’expressions simples. Les mauvaises options doivent avoir un certain degré de pertinence par rapport à la bonne réponse pour rendre la question difficile, mais sans mentionner l’entité de connexion entre les deux ($s_{1}$) ou d’autres détails.
@@
Question Generation Instruction Ich möchte eine deutsche Multi-Hop-Frage formulieren, basierend auf den gegebenen Wikidata-Wissens-Tripeln: ($s_{1}$, $r_{1}$, $o_{1}$) und ($o_{1}$, $r_{2}$, x2). Die Frage soll flüssig in deutscher Sprache gestellt werden und es ermöglichen, durch Abfrage von x0 die Information über x2 zu erhalten. Dabei muss die Frage die Beziehungen r1 und r2 explizit einbeziehen, ohne jedoch direkte oder indirekte Hinweise auf die Brückenentität x1 zu geben.
Option Generation Instruction Gegebene Informationen: ’question’ Zu den bekannten Fakten gehören: ($s_{1}$, $r_{1}$, $o_{1}$) und ($o_{1}$, $r_{2}$, x2). Generieren Sie auf der Grundlage dieser Informationen bitte eine korrekte Option A und drei falsche, aber vernünftig klingende Optionen B, C und D. Stellen Sie sicher, dass die Optionen als vollständige Sätze und nicht nur als einzelne Wörter oder Phrasen dargestellt werden. Die falschen Optionen sollten einen gewissen Grad an Relevanz für die richtige Antwort haben, um die Frage herausfordernd zu gestalten, ohne jedoch die verbindende Entität dazwischen zu erwähnen ($s_{1}$) oder andere Details.
@@
Appendix D Related Work
D.1 Knowledge Editing Datasets
D.2 Knowledge Editing Methods
Appendix E Multilingual Knowledge Editing Experiment